OpenStudy (anonymous):

Find all tangent lines to y = x^2 + 2x + 1 that pass through (0, -15)

OpenStudy (anonymous):

First find the derivative of that function

OpenStudy (anonymous):

Yeah, it's 2x + 2

OpenStudy (anonymous):

then plug in the x-coordinate value given by your point (this will be your slope of the derivative "tangent" line)

OpenStudy (anonymous):

The slope is 2

OpenStudy (anonymous):

Correct. Now using that, you can plug the coordinate and your new slope in the point slope formula \[y-y _{1}=m \left( x-x _{1} \right)\]

OpenStudy (anonymous):

y+ 15 = 2x

OpenStudy (anonymous):

If you get y to one side, then you'll notice that it looks like slope/intercept formula and that will be your tangent line

OpenStudy (anonymous):

Is that all the tangent lines?

OpenStudy (anonymous):

To my knowledge there can be multiple tangent lines to a function but only one per point (and you're only given one coordinate)

OpenStudy (freckles):

It doesn't say find the tangent line at (0,-15)

OpenStudy (anonymous):

Okay. It's just that I gave basically the same answer on my test for this question and got it wrong.

OpenStudy (anonymous):

It says it passes through, I thought it was the same thing

OpenStudy (freckles):

solve the following equation for a first: \[\frac{dy}{dx}|_{(x,y)=(a,f(a))}=\frac{f(a)-(-15)}{a-0}\]

OpenStudy (freckles):